How much Joules/cm^2 can a laser with wavelength 655nm and power 5mW deliver in 4 seconds? In how much time can it deliver 5 Joules? In how much time can it deliver 30 Joules?
1
Expert's answer
2013-05-02T11:27:23-0400
The power P (in our case 5 mW) is energy delivered per unit of time, that is P = E / t. The unit is 1 Watt = 1 Joule / second.

Without knowing the beam cross-section (thickness), we can't calculate the density of energy in Joules/cm^2, but we can calculate the integral energy in joules, that's what we will do. The wavelength of the radiation is irrelevant for these calculations.

a) in t = 4 seconds, the energy delivered is E = P*t = (5 mW) * (4 s) = 20 mJ.
b) to deliver E = 5 J, the time needed is t = E / P = (5 J) / (5 mW) = 1000 seconds =& 16.67 minutes.
c) to deliver E = 30 J, the time needed is t = E / P = (30 J) / (5 mW) = 6000 seconds = 100 minutes = 1.667 hours.